OpenStudy (anonymous):

3.2x+59.8=26.2-13.6x show work

OpenStudy (anonymous):

16.8x=-33.6 x=-2

OpenStudy (anonymous):

show work

OpenStudy (anonymous):

more detailed? just add and substractt lol

OpenStudy (anonymous):

3.2x+13.6x=26.2-59.8 16.8x=-33.6 x=-2

OpenStudy (anonymous):

3.2x + 59.8=26.2-13.6x +13.6x +13.6x ------------------------ 16.8x + 59.8=26.2 -59.8 -59.8 ------------------------ 16.8x=-33.6 ----- ---- 16.8 16.8 x=-2

OpenStudy (anonymous):

@Angel04